Metallic magnesium has a hexagonal close packed structure and has a density 1.74g cm-3. Assuming magnesium atoms to be spherical, calculate the radius of magnesium atom. (Atomic mass of Magnesium= 24.3).

Grade:upto college level

2 Answers

Balark Singhal
5 years ago
Magnesium is HCP
c/a =1.624
density =1.74g/cm3
For HCP n = 6
V = 3√3a2c/2 = 4.22a
Density = n×MW/Vc×Na
Na =6.02×1023atoms/mol
MW of magnesium = 24.3g/mol
V = 6×24.3/1.74×6.02×1023
a = 320pm
For HCP a =2R
R of magnesium= 160pm
 
sirikisai 143
one year ago
calculate the atimic radius of a hcp structure having c/a ratio of 1.5 and density of 1.25g/cm.atomic weight of atom is 24.31 g/mol